Determine if the sequence could be geometric. If so, identify the common ratio.
450, 150, 75, 25, …

answers:

a) yes; 0.5
b) yes; 1.5
c) yes; 2.5
d) no

Answer by htmentor(1343)

Determine if the sequence could be geometric. If so, identify the common ratio.
450, 150, 75, 25, …
There is no common ratio since:
a2/a1 = 150/450 = 1/3
a3/a2 = 75/150 = 1/2
Ans: d